JAWAHARLAL NEHRU TECHNOLOGICAL UNIVERSITY HYDERABAD

JAWAHARLAL NEHRU TECHNOLOGICAL UNIVERSITY HYDERABAD B.Tech II COMPUTER ORIENTED STATISTICAL METHODS Previous Question Paper (August/September – 2022 - R18)

B.Tech | II | II-I | R18 | August/September – 2022

Answer PaperPaper Code: 153AJMaximum Marks: 75Time: 3 Hours

Answer any five questions. All questions carry equal marks.

Main

115 marks
a

Three machines I, II and III produce 40%, 30% and 30% of the total number of items of a factory. The percentages of defective items of these machines are 4%, 2% and 3%. An item is selected at random and found to be defective. Find the probability that it is from:

i) Machine- I ii) Machine-II iii) Machine-III

8 marks

Answer

Let A1, A2, and A3 be the events that the item is produced by machines I, II, and III respectively, and let D be the event that the item is defective. The given probabilities are:

P(A1) = 0.40, P(A2) = 0.30, P(A3) = 0.30

P(D|A1) = 0.04, P(D|A2) = 0.02, P(D|A3) = 0.03

The total probability of selecting a defective item is:

P(D) = P(A1)P(D|A1) + P(A2)P(D|A2) + P(A3)P(D|A3)
P(D) = (0.40)(0.04) + (0.30)(0.02) + (0.30)(0.03) = 0.016 + 0.006 + 0.009 = 0.031

Using Bayes' Theorem, the posterior probabilities are:

i) P(A1|D) = (P(A1)P(D|A1)) / P(D) = 0.016 / 0.031 = 16/31 ≈ 0.516

ii) P(A2|D) = (P(A2)P(D|A2)) / P(D) = 0.006 / 0.031 = 6/31 ≈ 0.194

iii) P(A3|D) = (P(A3)P(D|A3)) / P(D) = 0.009 / 0.031 = 9/31 ≈ 0.290

b

A continuous Random variable has the p.d.f:

f(x) = e⁻ˣ for x ≥ 0, 0 otherwise

Determine: i) P(0 ≤ x ≤ 2) ii) The mean iii) Variance.

7 marks

Answer

i) P(0 ≤ x ≤ 2) = ∫02 e-x dx = [-e-x]02 = -e-2 - (-e0) = 1 - e-2 ≈ 1 - 0.1353 = 0.8647

ii) The mean (E[X]) is calculated as:

E[X] = ∫0∞ x e-x dx = 1! = 1

iii) To find the variance, first calculate E[X2]:

E[X2] = ∫0∞ x2 e-x dx = 2! = 2

Variance (V[X]) = E[X2] - (E[X])2 = 2 - (1)2 = 1

215 marks
a

There are three boxes. Box I contains 10 light bulbs (4 defective), Box II contains 6 light bulbs (1 defective), Box III contains 8 light bulbs (3 defective). A box is chosen at random and a bulb is selected. If it is defective find the probability that it is from:

i) Box- I ii) Box-II iii) Box-III

8 marks

Answer

Sign in to view the complete answer paper.

b

A continuous Random variable has the p.d.f:

f(x) = (1/2)(x+1) for -1 ≤ x ≤ 1, 0 otherwise

Determine: i) P(2 ≤ x ≤ 4) ii) The mean iii) Variance.

7 marks

Answer

Sign in to view the complete answer paper.

315 marks
a

Six cards are drawn from a pack of 52 cards. Getting a red card is a success. Find the probability of getting the success: i) At least once ii) 3 times

8 marks

Answer

Sign in to view the complete answer paper.

b

The probabilities of a Poisson variate taking the values 1 and 2 are equal. Find: i) μ ii) P(x≥1)

7 marks

Answer

Sign in to view the complete answer paper.

415 marks
a

Assume that 60% of the students passed an examination. Find the probability that among 12 students: i) Exactly 8 pass ii) At least 4 pass.

8 marks

Answer

Sign in to view the complete answer paper.

b

If the variance of a Poisson variate is 3. Find the probability that: i) P(x=0) ii) P(1 ≤ x < 4).

7 marks

Answer

Sign in to view the complete answer paper.

515 marks
a

In a test on electrical bulbs, it was found that the life was normally distributed with mean 2040 hours and S.D 40 hrs. Estimate the number of bulbs likely to burn for more than 2140 hours.

8 marks

Answer

Sign in to view the complete answer paper.

b

Two horses A, B were tested for time to run a track: Horse A: [28, 30, 32, 33, 33, 29, 34]; Horse B: [29, 30, 30, 24, 27, 29]. Test whether the two horses have the same running capacity at 95% level.

7 marks

Answer

Sign in to view the complete answer paper.

615 marks
a

If the masses of 300 students are normally distributed with mean 68 kg and S.D 3 kg. How many students have masses: i) Greater than 72 kg ii) Between 65 and 71 kg.

7 marks

Answer

Sign in to view the complete answer paper.

b

The following table gives the number of train accidents during the week. Test whether accidents are uniformly distributed at level 0.05.

DaysSunMonTuesWedThursFriSat
Accidents20181323261115
8 marks

Answer

Sign in to view the complete answer paper.

715 marks
a

A sample of 900 members has mean 3.4 cm and S.D 2.61 cm. Is this sample taken from a large population of mean 3.25 and S.D 2.61?

7 marks

Answer

Sign in to view the complete answer paper.

b

In city A, 20% of 900 school boys had a physical defect. In city B, 18.5% of 1600 school boys had the defect. Is the difference significant at 0.05 level?

8 marks

Answer

Sign in to view the complete answer paper.

8

The mobility of population from village, town, and city is given by the transition matrix:

To VillageTo TownTo City
From Village0.30.20.5
From Town0.30.50.2
From City0.10.40.5

Present population is 0.4, 0.3, 0.3 for village, town, and city. Find the proportions after two years and in the long run.

15 marks

Answer

Sign in to view the complete answer paper.

About this paper

This page contains the solved question paper with answers of COMPUTER ORIENTED STATISTICAL METHODS from JAWAHARLAL NEHRU TECHNOLOGICAL UNIVERSITY HYDERABAD for B.Tech CSE(SE) II II-I, August/September – 2022, R18.

For the question paper without answers, switch to the Question Paper tab.

Other papers in COMPUTER ORIENTED STATISTICAL METHODS

Loading Question Papers

Fetching data from database...